if f: R--->R such that f(x)f(y) – f(xy) = x+y where x,y belongs to R . find f(0)  where f(0) is not equal to 0 and hence find f(7).

Grade:11

1 Answers

Aman Sharma
7 years ago
Putting x and y =0,
f(0).f(0) – f(0) =0
f(0)- f(0)=0
f(0)[f(0)-1]=0
f(0) =1  or   f(0)=0
But f(0) = 0 is not equal to zero (Given)
 
Now Putting y =0,
f(0)f(x)-f(0)=x
f(x)=x+1
f(7)=8
f